OpenStudy (anonymous):

can someone walk me through how to solve this parabola??? x=1/2y^2 +3y+4

OpenStudy (anonymous):

By solve, what exactly do you mean?

OpenStudy (anonymous):

Solve for y?

OpenStudy (anonymous):

find an equation for it in standard form

OpenStudy (anonymous):

Find an equation? I'm confused.

OpenStudy (anonymous):

an equation... in standard form...

OpenStudy (anonymous):

Do you need it in the form x = a(x - h)^2 + k or x = ax^2 + bx + c?

OpenStudy (anonymous):

the first one. it's already in the second form

OpenStudy (anonymous):

i just keep messing up and need some help

OpenStudy (anonymous):

Oh. ok. THat's called general form where I come from and the second one is standard form. Ok, we're gonna do something called complete the square. First set x = 0 0=1/2y^2 +3y+4 Then, bring the constant to the other side. -4 = 1/2y^2 + 3y Now, remember to never complete the square unless the leading coefficient is 1. To get there, multiply the entire equation by 2. -8 = y^2 + 6y Now, you have to recognize that on the right side, it is starting to follow a perfect square trinomial. You have to complete the square by dividing 6 by 2 and then squaring it. -8 = y^2 + 6y + 9 Since this would change the equation, you also have to add a 9 on the other side to balance the equation. -8 + 9 = y^2 + 6y + 9 1 = y^2 + 6y + 9 Now it's starting to look like a perfect square trinomial right? You know how to factor it now. 1 = (y + 3)^2 To finish it, bring 1 back to the other side and set it equal to x again. x = (y + 3)^2 - 1

OpenStudy (anonymous):

Sorry that it's so long.

OpenStudy (anonymous):

thats ok... and i get what ur saying but my book is telling me a different answer

OpenStudy (anonymous):

oh. What's it say?

OpenStudy (anonymous):

What does your textbook say is the answer?

OpenStudy (anonymous):

x+1/2=1/2(y+3)^2

OpenStudy (anonymous):

Oh. I see what they did. Instead of multiplying by 2, they factored out 1/2. -4 = 1/2y^2 + 3y -4 = 1/2(y^2 + 6y + 9) -4 + 9/2 = 1/2(y^2 + 6y + 9) -4 + 9/2 = 1/2(y + 3)^2 1/2 = 1/2(y + 3)^2 Now, instead of bringing it to the other side, they just brought back x. x + 1/2 = 1/2(y + 3)^2

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

YES! thank u!!!

OpenStudy (anonymous):

You're welcome! :D

OpenStudy (anonymous):

and to find the directrix... can u possibly help with that?